4X² + y2 = 45 4(x-1) ² + y²= 25 Solve algebraically Check for non-Real solutions
Answer #1

expressed as 4277 y = 45 6. 4 (0-1) + y² = 25. (99) equation (11) can be 4 ( 22 - 2x+1) + y² = 25 => 472-84+4 +7= 25 => 422 +